# Gravitational wave probes of particle dark matter: a review

### v4

@@ VII.2 Dynamical friction on white dwarf binaries
Space-based GW detectors will observe the quasi-infinite inspiral of galactic white dwarf binaries and extreme-mass ratio inspirals of ordinary compact objects with much heavier BHs Bartolo et al. (2016); Babak et al. (2017). The long-durations of such signals permit the possibility of detecting a low-frequency modulation when passing through the ultralight dark matter soliton Wang and Zhong (2023); Brax et al. (2024); Blas et al. (2025). Ref. Brax et al. (2024) has found that the impact of dynamical friction on the phase evolution of GW signals from white dwarf binaries exceeds that of the low-frequency DM-induced modulation for masses above $10^{-21}$ eV, essentially providing two orders of magnitude of mass that could be probed in the future. Unfortunately, even at these masses, the density of DM would have to be at least $10^{4}$ times higher than in the solar system in order to observe any changes to the phase evolution GW. This is broadly consistent with previous work by Wang and Zhong (2023), which assumed an enhancement of the DM density by eight orders of magnitude relative to that on earth, and showed that DM with a mass of $\sim 10^{-23}$ eV could be detected with LISA. Furthermore, Ref. Blas et al. (2025) assumes a representative population of white dwarf binaries, a fraction of which are detectable in LISA, and concludes that DM masses above $\sim 10^{-23}$ eV will remain unconstrained by LISA. However, the mass regime $\sim[2\times 10^{-22},3\times 10^{-21}]$ eV could be constrained if DM has stronger, quadratic couplings to ordinary matter. Again, the possibility that DM couples to the standard model in some way is able to enhance the signal strength, as found in Sections II and IV for different types of DM.
@@ VII.3 Nanohertz modulations of continuous gravitational waves from neutron stars
In addition to modulating the millihertz frequency of GWs from binaries in space-based GW detectors, soliton ultralight dark matter could induce modulations in the rotational frequency of neutron stars in a way distinct from ordinary spinning down, deformed neutron stars, thus affecting CW signals and allowing the possibility of detecting this effect in future ground-based detectors Blas et al. (2025). These GWs are sourced from the rotational power of the neutron stars, thus spinning down the neutron stars as GWs are emitted Riles (2017). In ET and Cosmic Explorer, we expect to detect $\mathcal{O}(100)$ or more canonical neutron stars spinning down due to the emission of CWs Pagliaro et al. (2023), which could be modulated by ultralight dark matter solitons.
@@
Figure 23: Taken from Blas et al. (2025). Projected constraints on the linear (left) and quartic coupling parameters (right) for both LISA and ET from low-frequency soliton DM modulating (“heterodyning”) high-frequency carrier GW signals via changes to the moment of inertia of neutron stars, which alters CWs from deformed neutron stars (ET) and the GWs from inspiraling binary black holes (LISA). The curves for ET assume one or 200 detections of CWs from isolated neutron stars. The LISA curve is determined by the impact of soliton DM on GW frequencies from binary systems. The olive regions are excluded by the pulsar timing array analysis in Smarra et al. (2024), while blue and red are constrained away via Cassini tests of general relativity (violet) and on a stochastic GW background (red).
@@ VIII.1 Transmuted black holes in binaries
Non-annihilating WIMPs that interact with nucleons could lose energy through single or multiple scatterings off of celestial objects Bramante et al. (2017); Bell et al. (2021), eventually becoming gravitationally bound to them over time. After WIMPs repeatedly interact with ordinary nucleons, they become thermalized, forming a compact core with a higher density than that of the celestial object. If the density of the particles becomes sufficient inside the celestial object, gravitational collapse (in the case of bosons) or Chandrasekhar collapse (in the case of fermions) of the WIMP core will occur, thus resulting in a newborn BH that is extremely tiny with respect to the celestial object. After formation, the tiny BH could accrete surrounding DM, and if enough DM becomes bound to the object quickly enough compared to the Hawking evaporation timescale, the entire object be “swallowed” by the growing BH and become “transmuted” to what is known as a transmuted black hole (TBH) Dasgupta et al. (2021); Ray (2023); Bhattacharya et al. (2023, 2024); Bhattacharya (2024). Quantitatively, if the sum of the time to accrete new particles and the time for the WIMP core to become gravitationally unstable is smaller than the age of the universe, TBHs should form.
@@
TBHs forming from inspiraling neutron stars were first considered in Dasgupta et al. (2021). Since DM would induce neutron stars to transmute into BHs, from the observational perspective, there is no clear difference between electromagnetically silent neutron stars and low-mass BHs. Therefore, the non-detection of low-mass binary black holes by LIGO, Virgo and KAGRA can be used to constrain weak interaction cross-sections, since the progenitor neutron stars provide an extremely strong gravitational field to accrete and retain DM particles Bhattacharya et al. (2023); Bhattacharya (2024). In Fig. 24, we show the excluded bosonic DM mass/cross-section parameter space based on null search results from the LIGO, Virgo and KAGRA collaborations for sub-solar mass compact objects Abbott et al. (2023). Additionally, a projection for this constraint is shown for a 50x increase in the probable spacetime volume, which is reasonable for the future GW network using ET or Cosmic Explorer. While analyses of LIGO, Virgo and KAGRA data cannot yet constrain new parameter space for this model of DM, future searches should be able to surpass the pulsar constraint, which comes from the fact that the nearby Gyr-old pulsar PSR 0437-4715 has not actually collapsed into a BH in its lifetime McDermott et al. (2012); Garani et al. (2019).
@@
In contrast, stars, being less compact than neutron stars, could sustain DM cores with stronger interaction cross-sections, because of the weaker gravitational field around the star Bhattacharya et al. (2024). In Fig. 24, we show the potential excluded parameter space for null detection of TBHs arising from slowly inspiraling non-compact binaries that could be visible in space-based GW detectors. These binaries would be inspiraling for durations of order of the age of the universe, and thus exhibit CW signals in the millihertz regime. The results are parameterized in terms of $\alpha$, which denotes the fraction of stellar-mass binaries that form close enough to exist within the millihertz band. Essentially, $\alpha$ parametrizes our ignorance about how closely the population of sun-like stars form close to each other.
@@
Comparing Fig. 24 and Fig. 24, we note the vastly different interaction cross-section ranges from these two complementary GW probes of WIMP DM. For both low-mass BHs and sun-like stars, fermionic DM has also been constrained in similar ways as bosonic DM has, see Bhattacharya et al. (2023, 2024) for more details.
@@ VIII.2 GW probes of the galactic-center GeV excess
Annihilating WIMP DM and a collection of unresolved millisecond pulsars are two leading explanations for the unexpected excess of GeV gamma rays observed by the Fermi-LAT experiment coming from the Galactic Center Goodenough and Hooper (2009); Hooper and Goodenough (2011). At first, the GeV excess appeared spherically symmetric and well-fit by annihilating DM models Hooper and Linden (2011); Gordon and Macias (2013); Daylan et al. (2016); Calore et al. (2015a); Abazajian et al. (2014). But still, an astrophysical explanation of millisecond pulsars cannot yet be ruled out Abazajian (2011); Calore et al. (2015b); Yuan and Zhang (2014); Petrović et al. (2015); Ye and Fragione (2022).
GWs can be used to weigh in on this debate. If the millisecond pulsars are deformed, they will emit CWs as they rotate. The amplitude of these CWs is determined by both the GW frequency and the degree of deformation, called the ellipticity, which is defined as $\epsilon\equiv\frac{I_{\rm xx}-I_{\rm yy}}{I_{\rm zz}}$ Haskell et al. (2015), where $I_{\rm xx},I_{\rm yy}$ and $I_{\rm zz}$ are the principle moment of inertias along each axis.
One approach to constraining the millisecond pulsar hypothesis is to consider the bulk emission of a thousands or more sources of CWs that are superimposed, creating a stochastic GW background. Ref. Calore et al. (2019) modelled this stochastic GW background and projected constraints on the average ellipticity of this population as a function of the total number of millisecond pulsars in the Galactic Center. The authors showed that $\mathcal{O}(10^{4})$ millisecond pulsars with ellipticities of $\sim 10^{-7}$ could be detectable by current detectors and be consistent with the diffuse electromagnetic radiation coming from the Galactic Center, as shown in Fig. 25. However, they did not go so far as to link GW observations with Fermi-LAT ones.
Ref. Miller and Zhao (2023) took a different approach by analyzing the continuous gravitational wave (CGW) signals from individual neutron stars (NSs), specifically focusing on millisecond pulsars (MSPs). They explored two main scenarios for the deformation of these pulsars:
@@
    Magnetic Field-Induced Deformations: In the first scenario, they assumed that the deformations of the millisecond pulsars are sustained by the internal magnetic fields of the neutron stars, as proposed in previous work Lander et al. (2012); Mastrano and Melatos (2012); Lander (2014).
@@
Using these two models, Ref. Miller and Zhao (2023) then assumed that the ellipticity and frequency distributions of unresolved population of millisecond pulsars can be calcualted with knonw pulsars’ parameters Manchester et al. (2005). By integrating over these distributions with upper limits on ellipticities from an O3 LIGO search for isolated neutron stars Abbott et al. (2022b), they calculated the probability of detecting a CW from this population for both models of ellipticity noted above
In parallel, the number of millisecond pulsars in the Galactic Center can be inferred by modeling the electromagnetic emission associated with the GeV excess. This is done by fitting luminosity functions, which describe the expected number of millisecond pulsars needed to explain the GeV excess as a function of various model parameters Hooper and Linden (2016); Ploeg et al. (2020); Dinsmore and Slatyer (2022). By combining the predicted number of millisecond pulsars, derived from the luminosity functions, with the probability of detecting them via CWs, the authors were able to rule out certain luminosity function parameters. Specifically, they excluded models in which the number of millisecond pulsars would have been large enough to produce at least one detectable CW, thereby exploring the possible origins of the GeV excess.
@@ IX GW probes of DM spikes around inspiraling black holes
It was first recognized in Gondolo and Silk (1999) that BHs embedded in the DM halo of a galaxy could accrete DM adiabatically, potentially forming a high-density region around the BH known as a “spike.” Subsequent studies demonstrated that such DM spikes around supermassive BHs at centers of galaxies would have been disrupted by galactic mergers Merritt et al. (2002), if the BH formed off-center Ullio et al. (2001), or through interactions with stars in the surrounding stellar population Merritt (2004); Bertone and Merritt (2005). However, realistic profiles incorporating relativistic corrections have shown that these spikes can cause more DM to accumulate closer to the BHs in both Schwarzchild and Kerr metrics Sadeghian et al. (2013); Ferrer et al. (2017); Bertone (2024), which begs the question of whether these objects could survive galactic dynamics. Though it has been shown that heavy, particle-like DM would be dispersed in equal-mass merger events through N-body simulations Bertone et al. (2005), it is much less clear-cut for ultralight dark matter: multiple works Boudon et al. (2023); Zhong et al. (2023); Cardoso et al. (2022a, b); Vicente and Cardoso (2022); Annulli et al. (2020a, b); Brax et al. (2020); Ficarra (2021); Zhang et al. (2023b); Choudhary et al. (2021); Yang et al. (2018) suggest that such DM spikes could survive, even up to ten orbits before merger Aurrekoetxea et al. (2024). The extent to which DM spikes survive directly affects the amount of dynamical friction the binary will experience as it inspirals, thus altering the GWs from the inspiral.
@@
In light of these recent developments regarding the plausibility of forming DM spikes around supermassive BHs, Ref. Daniel et al. (2025) has forecast upper bounds on the DM densities around supermassive BHs using future LISA observations of these systems. These bounds provide new constraints on the DM spike formation and the nature of DM in the central regions of galaxies, offering a unique opportunity to test the high-density DM environment near supermassive BH binaries. Furthermore, Ref. Chan and Lee (2024) determines that the observed decay rate of the orbit can be completely explained by the existence of a DM spike around the supermassive BH in OJ 287, providing plausible evidence of the existence of DM spikes.
@@
Additionally, IMBHs can sustain DM spikes because their smaller masses and relatively longer timescales for mergers (with respect to those of supermassive BHs) prevent the spikes from being disrupted Zhao and Silk (2005); Kim et al. (2023). Similarly, recent studies suggest that both astrophysical BHs and PBHs could also host DM spikes if there is an absence of significant stellar interactions that could scatter DM particles, and if the gravitational influence of the IMBH persists over time, especially in environments where galactic mergers or other disruptive events have not destroyed the spike Kohri et al. (2014); Eroshenko (2016); Boucenna et al. (2018); Hertzberg et al. (2020); Bertone et al. (2025); Bertone (2024); Dosopoulou and Silk (2025). These findings expand the range of environments where DM spikes might form, suggesting that IMBHs could significantly influence DM distributions.
@@
For simplicity, despite the more realistic DM spike profiles recently discussed, the density of the DM spike is often approximated by the following equation Eda et al. (2013); Cole et al. (2023a, b):
@@
where $\rho_{\rm sp}$ is a normalization constant for the DM density of the spike, $r_{\rm min}$ is the minimum separation between the two objects, often taken to the radius at the innermost stable circular orbit, and $r_{\rm sp}$ is the spike radius. Here, $\rho_{\rm sp}=379\ M_{\odot}/\text{pc}^{3}$ and $r_{\rm sp}=0.33\text{pc}$ if the mass of the IMBH is $M_{\rm BH}=10^{3}M_{\odot}$ and the DM spike has a mass of $M_{\rm halo}=10^{6}M_{\odot}$. The power-law index $\gamma_{s}$ denotes how the concentration of DM particles falls off as one moves away from the host BH. Astrophysically speaking, measurements of $\gamma_{s}$ would allow us to glean important information regarding whether the host BH evolved from the DM halo and/or experienced mergers in the past Eda et al. (2013).

### v5

@@ VII.2 Dynamical friction on white dwarf binaries
Space-based GW detectors will observe the quasi-infinite inspiral of galactic white dwarf binaries and extreme-mass ratio inspirals of ordinary compact objects with much heavier BHs Bartolo:2016ami; Babak:2017tow. The long-durations of such signals permit the possibility of detecting a low-frequency modulation when passing through the ultralight dark matter soliton Wang:2023phr; Brax:2024yqh; Blas:2024duy. Ref. Brax:2024yqh has found that the impact of dynamical friction on the phase evolution of GW signals from white dwarf binaries exceeds that of the low-frequency DM-induced modulation for masses above $10^{-21}$ eV, essentially providing two orders of magnitude of mass that could be probed in the future. Unfortunately, even at these masses, the density of DM would have to be at least $10^{4}$ times higher than in the solar system in order to observe any changes to the phase evolution GW. This is broadly consistent with previous work by Wang:2023phr, which assumed an enhancement of the DM density by eight orders of magnitude relative to that on earth, and showed that DM with a mass of $\sim 10^{-23}$ eV could be detected with LISA. Furthermore, Ref. Blas:2024duy assumes a representative population of white dwarf binaries, a fraction of which are detectable in LISA, and concludes that DM masses above $\sim 10^{-23}$ eV will remain unconstrained by LISA. However, the mass regime $\sim[2\times 10^{-22},3\times 10^{-21}]$ eV could be constrained if DM has stronger, quadratic couplings to ordinary matter. Again, the possibility that DM couples to the standard model in some way is able to enhance the signal strength, as found in Sections II and IV for different types of DM.
@@ VII.3 Nanohertz modulations of continuous gravitational waves from neutron stars
In addition to modulating the millihertz frequency of GWs from binaries in space-based GW detectors, soliton ultralight dark matter could induce modulations in the rotational frequency of neutron stars in a way distinct from ordinary spinning down, deformed neutron stars, thus affecting CW signals and allowing the possibility of detecting this effect in future ground-based detectors Blas:2024duy. These GWs are sourced from the rotational power of the neutron stars, thus spinning down the neutron stars as GWs are emitted Riles:2017evm. In ET and Cosmic Explorer, we expect to detect $\mathcal{O}(100)$ or more canonical neutron stars spinning down due to the emission of CWs Pagliaro:2023bvi, which could be modulated by ultralight dark matter solitons.
@@
Figure 23: Taken from Blas:2024duy. Projected constraints on the linear (left) and quartic coupling parameters (right) for both LISA and ET from low-frequency soliton DM modulating (“heterodyning”) high-frequency carrier GW signals via changes to the moment of inertia of neutron stars, which alters CWs from deformed neutron stars (ET) and the GWs from inspiraling binary black holes (LISA). The curves for ET assume one or 200 detections of CWs from isolated neutron stars. The LISA curve is determined by the impact of soliton DM on GW frequencies from binary systems. The olive regions are excluded by the pulsar timing array analysis in Smarra:2024kvv, while blue and red are constrained away via Cassini tests of general relativity (violet) and on a stochastic GW background (red).
@@ VIII.1 Transmuted black holes in binaries
Non-annihilating WIMPs that interact with nucleons could lose energy through single or multiple scatterings off of celestial objects Bramante:2017xlb; Bell:2020lmm, eventually becoming gravitationally bound to them over time. After WIMPs repeatedly interact with ordinary nucleons, they become thermalized, forming a compact core with a higher density than that of the celestial object. If the density of the particles becomes sufficient inside the celestial object, gravitational collapse (in the case of bosons) or Chandrasekhar collapse (in the case of fermions) of the WIMP core will occur, thus resulting in a newborn BH that is extremely tiny with respect to the celestial object. After formation, the tiny BH could accrete surrounding DM, and if enough DM becomes bound to the object quickly enough compared to the Hawking evaporation timescale, the entire object be “swallowed” by the growing BH and become “transmuted” to what is known as a transmuted black hole (TBH) Dasgupta:2020mqg; Ray:2023auh; Bhattacharya:2023stq; Bhattacharya:2024pmp; Bhattacharya:2024cpm. Quantitatively, if the sum of the time to accrete new particles and the time for the WIMP core to become gravitationally unstable is smaller than the age of the universe, TBHs should form.
@@
TBHs forming from inspiraling neutron stars were first considered in Dasgupta:2020mqg. Since DM would induce neutron stars to transmute into BHs, from the observational perspective, there is no clear difference between electromagnetically silent neutron stars and low-mass BHs. Therefore, the non-detection of low-mass binary black holes by LIGO, Virgo and KAGRA can be used to constrain weak interaction cross-sections, since the progenitor neutron stars provide an extremely strong gravitational field to accrete and retain DM particles Bhattacharya:2023stq; Bhattacharya:2024cpm. In Fig. 24, we show the excluded bosonic DM mass/cross-section parameter space based on null search results from the LIGO, Virgo and KAGRA collaborations for sub-solar mass compact objects LIGOScientific:2022hai. Additionally, a projection for this constraint is shown for a 50x increase in the probable spacetime volume, which is reasonable for the future GW network using ET or Cosmic Explorer. While analyses of LIGO, Virgo and KAGRA data cannot yet constrain new parameter space for this model of DM, future searches should be able to surpass the pulsar constraint, which comes from the fact that the nearby Gyr-old pulsar PSR 0437-4715 has not actually collapsed into a BH in its lifetime McDermott:2011jp; Garani:2018kkd.
@@
In contrast, stars, being less compact than neutron stars, could sustain DM cores with stronger interaction cross-sections, because of the weaker gravitational field around the star Bhattacharya:2024pmp. In Fig. 24, we show the potential excluded parameter space for null detection of TBHs arising from slowly inspiraling non-compact binaries that could be visible in space-based GW detectors. These binaries would be inspiraling for durations of order of the age of the universe, and thus exhibit CW signals in the millihertz regime. The results are parameterized in terms of $\alpha$, which denotes the fraction of stellar-mass binaries that form close enough to exist within the millihertz band. Essentially, $\alpha$ parametrizes our ignorance about how closely the population of sun-like stars form close to each other.
@@
Comparing Fig. 24 and Fig. 24, we note the vastly different interaction cross-section ranges from these two complementary GW probes of WIMP DM. For both low-mass BHs and sun-like stars, fermionic DM has also been constrained in similar ways as bosonic DM has, see Bhattacharya:2023stq; Bhattacharya:2024pmp for more details.
@@ VIII.2 GW probes of the galactic-center GeV excess
Annihilating WIMP DM and a collection of unresolved millisecond pulsars are two leading explanations for the unexpected excess of GeV gamma rays observed by the Fermi-LAT experiment coming from the Galactic Center Goodenough:2009gk; Hooper:2010mq. At first, the GeV excess appeared spherically symmetric and well-fit by annihilating DM models Hooper:2011ti; Gordon:2013vta; Daylan:2014rsa; Calore:2014nla; Abazajian:2014fta. But still, an astrophysical explanation of millisecond pulsars cannot yet be ruled out Abazajian:2010zy; Calore:2014xka; Yuan:2014rca; Petrovic:2014xra; Ye:2022yxt.
GWs can be used to weigh in on this debate. If the millisecond pulsars are deformed, they will emit CWs as they rotate. The amplitude of these CWs is determined by both the GW frequency and the degree of deformation, called the ellipticity, which is defined as $\epsilon\equiv\frac{I_{\rm xx}-I_{\rm yy}}{I_{\rm zz}}$ Haskell:2015psa, where $I_{\rm xx},I_{\rm yy}$ and $I_{\rm zz}$ are the principle moment of inertias along each axis.
One approach to constraining the millisecond pulsar hypothesis is to consider the bulk emission of a thousands or more sources of CWs that are superimposed, creating a stochastic GW background. Ref. Calore:2018sbp modelled this stochastic GW background and projected constraints on the average ellipticity of this population as a function of the total number of millisecond pulsars in the Galactic Center. The authors showed that $\mathcal{O}(10^{4})$ millisecond pulsars with ellipticities of $\sim 10^{-7}$ could be detectable by current detectors and be consistent with the diffuse electromagnetic radiation coming from the Galactic Center, as shown in Fig. 25. However, they did not go so far as to link GW observations with Fermi-LAT ones.
Ref. Miller:2023qph took a different approach by analyzing the continuous gravitational wave (CGW) signals from individual neutron stars (NSs), specifically focusing on millisecond pulsars (MSPs). They explored two main scenarios for the deformation of these pulsars:
@@
    Magnetic Field-Induced Deformations: In the first scenario, they assumed that the deformations of the millisecond pulsars are sustained by the internal magnetic fields of the neutron stars, as proposed in previous work Lander:2011yr; Mastrano:2011aa; Lander:2013oea.
@@
Using these two models, Ref. Miller:2023qph then assumed that the ellipticity and frequency distributions of unresolved population of millisecond pulsars can be calcualted with knonw pulsars’ parameters Manchester:2004bp. By integrating over these distributions with upper limits on ellipticities from an O3 LIGO search for isolated neutron stars KAGRA:2022dwb, they calculated the probability of detecting a CW from this population for both models of ellipticity noted above
In parallel, the number of millisecond pulsars in the Galactic Center can be inferred by modeling the electromagnetic emission associated with the GeV excess. This is done by fitting luminosity functions, which describe the expected number of millisecond pulsars needed to explain the GeV excess as a function of various model parameters Hooper:2016rap; Ploeg:2020jeh; Dinsmore:2021nip. By combining the predicted number of millisecond pulsars, derived from the luminosity functions, with the probability of detecting them via CWs, the authors were able to rule out certain luminosity function parameters. Specifically, they excluded models in which the number of millisecond pulsars would have been large enough to produce at least one detectable CW, thereby exploring the possible origins of the GeV excess.
@@ IX GW probes of DM spikes around inspiraling black holes
It was first recognized in Gondolo:1999ef that BHs embedded in the DM halo of a galaxy could accrete DM adiabatically, potentially forming a high-density region around the BH known as a “spike.” Subsequent studies demonstrated that such DM spikes around supermassive BHs at centers of galaxies would have been disrupted by galactic mergers Merritt:2002vj, if the BH formed off-center Ullio:2001fb, or through interactions with stars in the surrounding stellar population Merritt:2003qk; Bertone:2005hw. However, realistic profiles incorporating relativistic corrections have shown that these spikes can cause more DM to accumulate closer to the BHs in both Schwarzchild and Kerr metrics Sadeghian:2013laa; Ferrer:2017xwm; Bertone:2024rxe, which begs the question of whether these objects could survive galactic dynamics. Though it has been shown that heavy, particle-like DM would be dispersed in equal-mass merger events through N-body simulations Bertone:2004pz, it is much less clear-cut for ultralight dark matter: multiple works Boudon:2023qbu; Zhong:2023xll; Cardoso:2022nzc; Cardoso:2022vpj; Vicente:2022ivh; Annulli:2020lyc; Annulli:2020ilw; Brax:2019npi; Ficarra:2021qeh; Zhang:2022rex; Choudhary:2020pxy; Yang:2017lpm suggest that such DM spikes could survive, even up to ten orbits before merger Aurrekoetxea:2024cqd. The extent to which DM spikes survive directly affects the amount of dynamical friction the binary will experience as it inspirals, thus altering the GWs from the inspiral.
@@
In light of these recent developments regarding the plausibility of forming DM spikes around supermassive BHs, Ref. Daniel:2025mna has forecast upper bounds on the DM densities around supermassive BHs using future LISA observations of these systems. These bounds provide new constraints on the DM spike formation and the nature of DM in the central regions of galaxies, offering a unique opportunity to test the high-density DM environment near supermassive BH binaries. Furthermore, Ref. Chan:2024yht determines that the observed decay rate of the orbit can be completely explained by the existence of a DM spike around the supermassive BH in OJ 287, providing plausible evidence of the existence of DM spikes.
@@
Additionally, IMBHs can sustain DM spikes because their smaller masses and relatively longer timescales for mergers (with respect to those of supermassive BHs) prevent the spikes from being disrupted Zhao:2005zr; Kim:2022mdj. Similarly, recent studies suggest that both astrophysical BHs and PBHs could also host DM spikes if there is an absence of significant stellar interactions that could scatter DM particles, and if the gravitational influence of the IMBH persists over time, especially in environments where galactic mergers or other disruptive events have not destroyed the spike Kohri:2014lza; Eroshenko:2016yve; Boucenna:2017ghj; Hertzberg:2019exb; Bertone:2024wbn; Bertone:2024rxe; Dosopoulou:2025jth. These findings expand the range of environments where DM spikes might form, suggesting that IMBHs could significantly influence DM distributions.
@@
For simplicity, despite the more realistic DM spike profiles recently discussed, the density of the DM spike is often approximated by the following equation Eda:2013gg; Cole:2022ucw; Cole:2022yzw:
@@
where $\rho_{\rm sp}$ is a normalization constant for the DM density of the spike, $r_{\rm min}$ is the minimum separation between the two objects, often taken to the radius at the innermost stable circular orbit, and $r_{\rm sp}$ is the spike radius. Here, $\rho_{\rm sp}=379\ M_{\odot}/\text{pc}^{3}$ and $r_{\rm sp}=0.33\text{pc}$ if the mass of the IMBH is $M_{\rm BH}=10^{3}M_{\odot}$ and the DM spike has a mass of $M_{\rm halo}=10^{6}M_{\odot}$. The power-law index $\gamma_{s}$ denotes how the concentration of DM particles falls off as one moves away from the host BH. Astrophysically speaking, measurements of $\gamma_{s}$ would allow us to glean important information regarding whether the host BH evolved from the DM halo and/or experienced mergers in the past Eda:2013gg.